Suppose $X$ has Poisson distribution such that $3 P(X=1)=2 P(X=2)$ then $P(X>0)$ is:
Answer & explanation
Correct answer: option 3
The correct answer is Option (3) → $1-e^{-3}$
$X\sim \text{Poisson}(\lambda)$
$P(X=x)=\frac{e^{-\lambda}\lambda^x}{x!}$
$P(X=1)=e^{-\lambda}\lambda$
$P(X=2)=\frac{e^{-\lambda}\lambda^2}{2}$
$3P(X=1)=2P(X=2)$
$3(e^{-\lambda}\lambda)=2\left(\frac{e^{-\lambda}\lambda^2}{2}\right)$
$3\lambda=\lambda^2$
$\lambda=3$
$P(X>0)=1-P(X=0)$
$P(X=0)=e^{-3}$
$P(X>0)=1-e^{-3}$
The probability is $1-e^{-3}$.